# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 6.7 Proofs of Basic Facts for Dirichlet Form and Variance
6.7.1 Proof of Lemma 6.3
@@
$=\sum_{\begin{subarray}{c}\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}:\\ \sigma_{1}\cap\sigma_{2}=\eta\in\mathcal{P}_{j}\end{subarray}}\frac{\pi_{i}(\sigma_{1})\pi_{i}(\sigma_{2})}{\pi_{j}(\eta)\times(j+1)\times\dots\times n)^{2}}f^{(i)}(\sigma_{1})f^{(i)}(\sigma_{2})$ (41)
@@
$=:\mathrm{Var}_{\pi_{i}}f^{(k)}-\mathrm{Var}_{\pi_{j}}f^{(j)}$ by Eqs. 38 and 42
where the second line follows from the reversibility condition $\pi_{i}(\sigma)\mathrm{P}_{i}^{\vee}(\sigma,\tau)=\pi_{i}(\tau)\mathrm{P}_{i}^{\vee}(\tau,\sigma)$ and the fact that $\sum_{\tau\in\mathcal{P}_{i}}\mathrm{P}_{i}^{\vee}(\sigma,\tau)=1$. This completes the proof of the lemma.
6.7.2 Proof of Lemma 6.4
We will use the following identity in the proof of Lemma 6.4.
@@
And for the variance at level $k-1$ we have:
@@
for every $\mbox{for every $\eta\in\mathcal{P}_{k-1}$, }f^{(k-1)}(\eta)=\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)$ (46)
for every $\mbox{for every $\sigma\in\mathcal{P}_{k+1}$, }\sum_{\begin{subarray}{c}\eta\in\mathcal{P}_{k-1}:\\ \eta\subset\sigma\end{subarray}}\pi_{k-1}(\eta)\pi_{\eta,2}(\sigma\setminus\eta)=\pi_{k+1}(\sigma)$ (47)
@@
If we used the bound $1/(1-\gamma_{k-1}/2)$ in (49) instead of $2\gamma_{k-1}$ then the statement of Lemma 6.4 has $\gamma_{k-1}/(2-\gamma_{k-1})$ in place of $(2\gamma_{k-1}-1)$; this matches the statement of [CLV21, Fact A.8 and Theorem A.9]. ∎
7 Optimal Mixing Time via Entropy Decay
In the previous section we established Theorem 2 which says that spectral independence implies optimal $O(n)$ relaxation time of the Glauber dynamics. Theorem 3 is the stronger result, which says that spectral independence and marginal boundedness implies optimal $O(n\log{n})$ mixing time. Both Theorems 2 and 3 were proved by Chen, Liu, and Vigoda [CLV21], and our proofs of these theorems follow their proof approach. To obtain Theorem 3, we follow the same proof approach as in the proof of Theorem 2 with the general objective of replacing $\mathrm{Var}$ by $\mathrm{Ent}$.
There are two main sticking points with replacing $\mathrm{Var}$ by $\mathrm{Ent}$. The first is that the proof of Theorem 2 uses the spectral gap of the Glauber dynamics. Recall the formulation of the spectral gap in the Poincare inequality, see Eq. 2, where the spectral gap $\gamma=\min_{f}\mathcal{E}(f)/\mathrm{Var}(f)$. Moreover, for the down-up chain $\mathrm{P}_{i,j}^{\vee}$ one can replace the $\mathcal{E}_{\mathrm{P}_{i,j}^{\vee}}(f)$ by $\mathrm{Var}_{\pi_{i}}(f)-\mathrm{Var}_{\pi_{j}}(f)$, see Lemma 6.3. Notice that in Section 6 for the proof of Theorem 2 we only considered the down-up chain and hence we can rephrase the proof only in terms of $\mathrm{Var}(f)$, without use of $\mathcal{E}(f)$. Therefore we can aim to replace all occurrences of $\mathrm{Var}(f)$ with $\mathrm{Ent}(f)$ in order to establish entropy decay instead of variance decay.
The only technical obstacle in replacing $\mathrm{Var}$ by $\mathrm{Ent}$ is in the application of Theorem 1 in the proof of Theorem 2 in Section 6.5. Theorem 1 established a lower bound on the spectral gap of the Glauber dynamics but it does not establish a bound on its entropy decay. To get a corresponding bound on entropy decay we need to use an additional property known as marginal boundedness, see Definition 1.4. Using $b$-marginal boundedness, [CLV21, Lemma 4.2] proved a corresponding bound on $\mathrm{Ent}_{C}(f)$ in terms of $\sum_{v\in C}\mu(\mathrm{Ent}_{v}(f))$ which depends on $b$ and $|C|$, and this bound can be used in the analogous step for establishing entropy decay, see the proof of Lemma 2.3 in [CLV21].
The analog of the proof of Theorem 2 in Section 6.5, as outlined above with $\mathrm{Ent}$ in place of $\mathrm{Var}$, yields entropy tensorization of the Gibbs distribution as formulated in the following definitions.
@@ 7 Optimal Mixing Time via Entropy Decay
Finally, we can define entropy tensorization.
@@
A distribution $\mu$ with support $\Omega\subseteq\{0,1\}^{V}$ satisfies approximate tensorisation of entropy with constant $C>0$, if for all $f:\Omega\to\mathbb{R}_{\geq 0}$ we have that
@@
Note, the corresponding notion of variance tensorization is equivalent to the spectral gap of the (heat-bath) Glauber dynamics as the RHS in the above is equivalent to the Dirichlet form, see Eq. 27.
@@
$T_{\text{mix}}\leq Cn\log(\log(1/\mu^{*})).$ (50)
This outlines the approach of [CLV21] to obtain optimal mixing time bounds.
Further improvements are obtained in [CFYZ22, CE22] which bound the mixing time as $\leq C(\eta)n\log{n}$ instead of $C(\eta,\Delta)$ based on a stronger form of spectral independence, and hence their results potentially extend to unbounded maximum degree $\Delta$.
@@ 8 Matroids
We now utilize the spectral independence approach to establish fast mixing of a Markov chain for generating a random basis of a matroid. The main result for matroids is work of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19]. In fact, the work of [ALOV19] for matroids inspired the idea of spectral independence as we presented earlier and introduced in [ALO20], however we present these results in the opposite order.
The general approach to prove fast mixing of the chain on bases of a matroid is to apply the Random Walk Theorem (Theorem 4). To bound the spectral gap of the local walks we will use a result of Oppenheim [Opp18] known as the Trickle-Down Theorem. The Trickle-Down Theorem bounds the spectral gap of the local walk $Q_{S}$ for a pinning $S$ by induction on $|S|$. The base case will be a bound on the gap of the local walk $Q_{S}$ when $k:=|S|=n-2$ which will correspond to rank-2 matroids; this is a significantly easier case to analyze as it corresponds to an unweighted walk. We will then use the Trickle-Down Theorem to deduce a bound on the spectral gap of $Q_{S}$ for pinnings of size $k-1$ and hence any pinning $S$ by induction.
@@ 8.1 Matroids: Definitions
Exercise 8.3.
Let $\mathcal{F}$ be a set of bases of a matroid $\mathcal{M}$. Let $\mathcal{F}^{*}$ be the set of complements, that is, *=$\mathcal{F}^{*}=\{[n]-B:B\in\mathcal{F}\}$. Show that $\mathcal{F}^{*}$ are the bases of a matroid $\mathcal{M}^{*}$ (called the dual of the matroid $\mathcal{M}$).
Exercise 8.4.
Let $\mathcal{I}$ be a set of independent sets of a matroid $M=(E,\mathcal{I})$. Let $S\subseteq[n]$. Let $\mathcal{I}^{\prime\prime}$ be the set of independent sets contained in $S$, that is, $\{T:T\in\mathcal{I}\ \mbox{and}\ T\subseteq S\}$. Show that ′′$\mathcal{M}^{\prime\prime}=(E,\mathcal{I}^{\prime\prime})$ is a matroid (called the restriction of the matroid $\mathcal{M}$; denoted by $\mathcal{M}|S$).
Exercise 8.5.
@@ 8.2 Matroids: Examples
Let $G=(V,E)$ be a connected graph. Let =$\mathcal{I}=\{S\subset E:(V,S)\mbox{ is acyclic}\}$ be the collection of all forests of $G$. The collection $\mathcal{M}=(E,\mathcal{I})$ is a matroid. The bases of $\mathcal{M}$ are maximal acyclic subgraphs. Since $G$ is assumed to be connected then the bases of $\mathcal{M}$ are the spanning trees of $G$ since $G$, and hence this is also called the spanning tree matroid.
@@
=$\mathcal{I}=\{S\subseteq V:S\text{ is linear independent}\}.$
@@
8.3 Bases Exchange Walk
Our central question is, for a given matroid $\mathcal{M}=(E,\mathcal{I})$, can we generate a basis, uniformly at random from the set $\mathcal{F}$ of all bases, in time polynomial in $n=|E|$. We are given the matroid in implicit form, so we can efficiently check whether a given subset $S\subset E$ is independent or dependent. We will use the following natural Markov chain known as the bases exchange walk to generate a random bases.
The bases exchange walk is the following Markov chain $(B_{t})$ on the set of bases of a given matroid $\mathcal{M}$. Let $\mathcal{F}$ denote the collection of bases of a matroid $M$. From $B_{t}\in\mathcal{F}$, the transitions $B_{t}\rightarrow B_{t+1}$ are defined as follows:
@@ 8.4 Main Theorem: Fast mixing of Bases-Exchange Walk
Theorem 6 ([ALOV19]).
For any matroid $M$, the spectral gap of the bases-exchange walk satisfies $\geq\frac{1}{r(M)}$, where $r(M)$ is the rank of the matroid.
Using the upper bound on the mixing time in terms of the spectral gap (see Eq. 4) we have the following.
Corollary 8.8.
The mixing time of the bases-exchange walk is $O(r(M)\log(|\mathcal{F}|))=O(r(M)^{2}\log n)$.
By proving entropy contraction, [CGM21] improves the mixing time bound to $O(r(M)\log{\log(|\mathcal{F}|)})=O(r(M)\log r(M)+r(M)\log{\log{n}})$; this is presented in Heng Guo’s first lecture. Furthermore, the $\log\log n$ term was subsequently removed in [${}^{+}$21], achieving $O(r(M)\log r(M))$ mixing time. This is tight because we need $\Omega(r(M)\log r(M))$ time to replace every element due to coupon collector.

### v4

@@ 6.7 Proofs of Basic Facts for Dirichlet Form and Variance
Here we provide the missing proofs of Lemmas 6.4 and 6.5. We begin with the proof of Lemma 6.4.
@@
$=\sum_{\begin{subarray}{c}\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}:\\ \sigma_{1}\cap\sigma_{2}=\eta\in\mathcal{P}_{j}\end{subarray}}\frac{\pi_{i}(\sigma_{1})\pi_{i}(\sigma_{2})}{\pi_{j}(\eta)\times(j+1)\times\dots\times n)^{2}}f^{(i)}(\sigma_{1})f^{(i)}(\sigma_{2})$
@@
$=:\mathrm{Var}_{\pi_{i}}f^{(k)}-\mathrm{Var}_{\pi_{j}}f^{(j)}$ by Eqs. 63 and 64
where the second line follows from the reversibility condition $\pi_{i}(\sigma)\mathrm{P}_{i}^{\vee}(\sigma,\tau)=\pi_{i}(\tau)\mathrm{P}_{i}^{\vee}(\tau,\sigma)$ and the fact that $\sum_{\tau\in\mathcal{P}_{i}}\mathrm{P}_{i}^{\vee}(\sigma,\tau)=1$. This completes the proof of the lemma. ∎
We now prove Lemma 6.5. We will use the following identity in the proof of Lemma 6.5.
@@
And for the variance at level k1k-1 we have:
@@
$\mbox{for every $\eta\in\mathcal{P}_{k-1}$, }f^{(k-1)}(\eta)=\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)$ (68)
$\mbox{for every $\sigma\in\mathcal{P}_{k+1}$, }\sum_{\begin{subarray}{c}\eta\in\mathcal{P}_{k-1}:\\ \eta\subset\sigma\end{subarray}}\pi_{k-1}(\eta)\pi_{\eta,2}(\sigma\setminus\eta)=\pi_{k+1}(\sigma)$ (69)
@@
If we used the bound $1/(1-\gamma_{k-1}/2)$ in (71) instead of $2\gamma_{k-1}$ then the statement of Lemma 6.5 has $\gamma_{k-1}/(2-\gamma_{k-1})$ in place of $(2\gamma_{k-1}-1)$; this matches the statement of [CLV21a, Fact A.8 and Theorem A.9]. ∎
6.8 Optimal Mixing Time via Entropy Decay
In the previous section we established Theorem 1.5 which says that spectral independence implies optimal $O(n)$ relaxation time of the Glauber dynamics. In contrast, Theorem 1.8 says that spectral independence and marginal boundedness implies optimal $O(n\log{n})$ mixing time. Both Theorems 1.5 and 1.8 were proved by Chen, Liu, and Vigoda [CLV21a], and our proofs of these theorems follow their proof approach. To obtain Theorem 1.8, we follow the same proof approach as in the proof of Theorem 1.5 with the general objective of replacing the variance functional $\mathrm{Var}$ by the entropy functional $\mathrm{Ent}$.
@@ 6.8 Optimal Mixing Time via Entropy Decay
We utilize entropy tensorization, which is the analog of approximate tensorization of variance, see Definition 3.4.
@@
A distribution $\mu$ with support $\Omega\subseteq\{0,1\}^{V}$ satisfies approximate tensorization of entropy with constant $C>0$, if for all $f:\Omega\to\mathbb{R}_{\geq 0}$ we have that
@@
Note, the corresponding notion of variance tensorization is equivalent to the spectral gap of the (heat-bath) Glauber dynamics as the RHS in the above is equivalent to the Dirichlet form, see Corollary 3.5.
@@
$T_{\mathrm{mix}}\leq Cn\log(\log(1/\mu^{*})).$ (72)
As suggested above, to establish approximate tensorization of entropy we aim to replace $\mathrm{Var}$ by $\mathrm{Ent}$ in the proof of Theorem 1.5. The first is that the proof of Theorem 1.5 uses the spectral gap of the Glauber dynamics. Recall the formulation of the spectral gap in the Poincaré inequality, see Eq. 15, where the spectral gap $\gamma=\min_{f}\mathcal{E}(f)/\mathrm{Var}(f)$. Moreover, for the down-up chain $\mathrm{P}_{i,j}^{\vee}$ one can replace the $\mathcal{E}_{\mathrm{P}_{i,j}^{\vee}}(f)$ by $\mathrm{Var}_{\pi_{i}}(f)-\mathrm{Var}_{\pi_{j}}(f)$, see Lemma 6.4. Notice that in Section 6 for the proof of Theorem 1.5 we only considered the down-up chain and hence we can rephrase the proof only in terms of $\mathrm{Var}(f)$, without use of $\mathcal{E}(f)$. Therefore we can aim to replace all occurrences of $\mathrm{Var}(f)$ with $\mathrm{Ent}(f)$ in order to establish entropy decay instead of variance decay.
The only technical obstacle in replacing $\mathrm{Var}$ by $\mathrm{Ent}$ is in the application of Theorem 1.4 in the proof of Theorem 1.5 in Section 6.5; in particular, see Eq. 54. Theorem 1.4 established a lower bound on the spectral gap of the Glauber dynamics but it does not establish a bound on its entropy decay. To get an analgous bound on entropy decay we need to use an additional property known as marginal boundedness, see Definition 1.7. Using $b$-marginal boundedness, [CLV21a, Lemma 4.2] proved a corresponding bound on $\mathrm{Ent}_{T}(f)$ in terms of $\sum_{v\in T}\mu(\mathrm{Ent}_{v}(f))$ which depends on $b$ and $|T|$:
$\mathrm{Ent}_{T}(F|\tau)\leq C^{\prime}\sum_{v\in T}\mu_{T}^{\tau}(\mathrm{Ent}_{v}[F|\tau]).$ (73)
In [CLV21a, Lemma 4.2] they established $C^{\prime}=C^{\prime}(k,b)=\frac{3k^{2}\log(1/b)}{2b^{2k+2}}$ where $k=|T|$ (note $b$ typically depends on $\Delta$, and also depends on $\lambda$ or $\eta$). Below, we provide a simple proof that $C^{\prime}=C^{\prime}(|T|,b,\Delta,\eta)=2C(\eta,\Delta)k\log(1/b)$ where $C(\eta,\Delta)$ is the constant from Theorem 1.5. The bound in Eq. 73 (with either constant $C^{\prime}$) can be used in the analogous step of Eq. 54 to instead establish entropy decay, and then conclude approximate tensorization of entropy of the Gibbs distribution (see Definition 6.8 above) which implies an optimal $O(n\log{n})$ bound on the mixing time, via Eq. 72.
This outlines the approach of [CLV21a] to obtain optimal mixing time bounds. Further improvements are obtained in [CFYZ22, CE22]; these works upper bound the mixing time by $C(\eta)n\log{n}$ instead of $C(\eta,\Delta)n\log{n}$ based on a stronger form of spectral independence. The work of Chen, Feng, Yin, and Zhang [CFYZ22] introduce the field dynamics which is presented in Section 10. In contrast, the work of Chen and Eldan [CE22] apply the framework of stochastic localization.
Proof of Eq. 73.
Consider the graph $G=(V,E)$ where $T\subseteq V$. Let $k=|T|$. We are running the Glauber dynamics on $T$ where for every $v\notin T$ the configuration at $v$ is fixed to $\tau(v)$. Let $\mu=\mu_{\tau}$ denote the Gibbs distribution on $T$ with fixed configuration $\tau$ outside $T$.
From Diaconis and Saloff-Coste [DS96] (see the proof of Corollary A.4), we know the following inequality:
$\mathrm{Ent}_{\mu}(f)\leq\left[\frac{\log(1/\mu^{*})-1}{1-2\mu^{*}}\right]\mathrm{Var}_{\mu}(\sqrt{f}),$ (74)
where $\mu^{*}=\min_{\sigma\in\Omega}\mu(\sigma)$. From Theorem 1.5 and Corollary 3.5, we have approximation tensorization of variance:
$\mathrm{Var}_{\mu}(f)\leq C(\eta,\Delta)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)].$ (75)
We will use the following elementary fact (e.g., see [DS96]), which we will prove momentarily:
$\mathrm{Var}_{v}(\sqrt{f})\leq\mathrm{Ent}_{v}(f).$ (76)
Therefore, by combining Eq. 74 and Eq. 75, and since $\mu^{*}\geq b^{k}$ we have:
$\mathrm{Ent}_{\mu}(f)$ $\leq C(\eta,\Delta)\left[\frac{\log(1/\mu^{*})-1}{1-2\mu^{*}}\right]\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(\sqrt{f})]$ by Eq. 74 (77)
$\leq 2C(\eta,\Delta)k\log(1/b)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(\sqrt{f})]$ by Eq. 75 (78)
$\leq 2C(\eta,\Delta)k\log(1/b)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Ent}_{v}(f)]$ by Eq. 76. (79)
This completes the proof of Eq. 73 with constant $C^{\prime}=2C(\eta,\Delta)k\log(1/b)$.
For completeness we now prove Eq. 76. Without loss of generality, we assume $f$ satisfies $\mathbb{E}_{\mu}[f]=1$. We then have the following:
$\mathrm{Ent}_{\mu}(f)$ $=\mathbb{E}_{\mu}[f\log(f)]$
$=2\mathbb{E}_{\mu}[f\log(\sqrt{f})]$
$\geq 2\mathbb{E}_{\mu}\left[f\left(1-\frac{1}{\sqrt{f}}\right)\right]$ (80)
$=2(1-\mathbb{E}_{\mu}[\sqrt{f}])$
$\geq 1-(\mathbb{E}_{\mu}[\sqrt{f}])^{2}$ (81)
$=\mathrm{Var}_{\mu}(\sqrt{f}),$
where Eq. 80 follows from the fact that $\log(x)=\log(1+x-1)\leq x-1$ and hence $\log(1/x)\geq 1-x$, and Eq. 81 follows from $2(1-x)\geq 1-x^{2}$. ∎
@@ 7 Matroids
We now utilize the spectral independence approach to establish fast mixing of a Markov chain for generating a random basis of a matroid. The main result for matroids is the work of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19]. In fact, the work of [ALOV19] for matroids inspired the idea of spectral independence as we presented earlier and introduced in [ALO20], however we present these results in the opposite order.
The general approach to prove fast mixing of the chain on bases of a matroid is to apply the Random Walk Theorem (Theorem 4.1). To bound the spectral gap of the local walks we will use a result of Oppenheim [Opp18] known as the Trickle-Down Theorem. The Trickle-Down Theorem bounds the spectral gap of the local walk $Q_{S}$ for a pinning $S$ by induction on $|S|$. The base case will be a bound on the gap of the local walk $Q_{S}$ when k:=|S|=n2k:=|S|=n-2 which will correspond to rank-2 matroids; this is a significantly easier case to analyze as it corresponds to an unweighted walk. We will then use the Trickle-Down Theorem to deduce a bound on the spectral gap of $Q_{S}$ for pinnings of size k1k-1 and hence any pinning $S$ by induction.
@@ 7.1 Matroids: Definitions
Exercise 7.3.
Let $\mathcal{F}$ be a set of bases of a matroid $\mathcal{M}$. Let $\mathcal{F}^{*}$ be the set of complements, that is, $\mathcal{F}^{*}=\{[n]-B:B\in\mathcal{F}\}$. Show that $\mathcal{F}^{*}$ are the bases of a matroid $\mathcal{M}^{*}$ (called the dual of the matroid $\mathcal{M}$).
Exercise 7.4.
Let $\mathcal{I}$ be a set of independent sets of a matroid $M=(E,\mathcal{I})$. Let $S\subseteq[n]$. Let $\mathcal{I}^{\prime\prime}$ be the set of independent sets contained in $S$, that is, $\{T:T\in\mathcal{I}\ \mbox{and}\ T\subseteq S\}$. Show that $\mathcal{M}^{\prime\prime}=(E,\mathcal{I}^{\prime\prime})$ is a matroid (called the restriction of the matroid $\mathcal{M}$; denoted by $\mathcal{M}|S$).
Exercise 7.5.
@@ 7.2 Matroids: Examples
Let $G=(V,E)$ be a connected graph. Let $\mathcal{I}=\{S\subset E:(V,S)\mbox{ is acyclic}\}$ be the collection of all forests of $G$. The collection $\mathcal{M}=(E,\mathcal{I})$ is a matroid. The bases of $\mathcal{M}$ are maximal acyclic subgraphs. Since $G$ is assumed to be connected then the bases of $\mathcal{M}$ are the spanning trees of $G$ since $G$, and hence this is also called the spanning tree matroid.
@@
$\mathcal{I}=\{S\subseteq V:S\text{ is linear independent}\}.$
@@
7.3 Bases-exchange Walk
Our central question is, for a given matroid $\mathcal{M}=(E,\mathcal{I})$, can we generate a basis, uniformly at random from the set $\mathcal{F}$ of all bases, in time polynomial in $n=|E|$. We are given the matroid in implicit form, so we can efficiently check whether a given subset $S\subset E$ is independent or dependent. We will use the following natural Markov chain known as the bases-exchange walk to generate a random bases.
The bases-exchange walk is the following Markov chain $(B_{t})$ on the set of bases of a given matroid $\mathcal{M}$. Let $\mathcal{F}$ denote the collection of bases of a matroid $M$. From $B_{t}\in\mathcal{F}$, the transitions $B_{t}\rightarrow B_{t+1}$ are defined as follows:
@@ 7.4 Main Theorem: Fast Mixing of Bases-Exchange Walk
See 1.15
By proving entropy contraction, [CGM21] improves the mixing time bound to $O(r(M)\log{\log(|\mathcal{F}|)})=O(r(M)\log r(M)+r(M)\log{\log{n}})$. Furthermore, the $\log\log n$ term was subsequently removed in [ALO+21], achieving $O(r(M)\log r(M))$ mixing time. This is tight because we need $\Omega(r(M)\log r(M))$ time to replace every element due to coupon collector.